OpenStudy (anonymous):

This year your brother Jack will be 2 years from being twice as old as your sister Jen. The sum of Jack's age and three times Jens age is 66. How old is Jen?

OpenStudy (anonymous):

could you type that equation out for me?

OpenStudy (anonymous):

b+2=1/2 *(s) b+3s=66

OpenStudy (anonymous):

i guess use the substitution method to solve

OpenStudy (anonymous):

i got b = 54/7, s = 136/7 b = 7.714, s = 19.428

OpenStudy (anonymous):

I think i might have made a mistake nvrmnd

OpenStudy (amistre64):

\[2b\pm 2 = s\]

OpenStudy (amistre64):

er.. \[2(b\pm2)=s\] sounds better to me

OpenStudy (amistre64):

and since\[b+3s=66\]subbing in is good\[b+3(2(b\pm2)) = 66\]

OpenStudy (amistre64):

i hate words in math ....

OpenStudy (amistre64):

\(b \pm 2\) = 2s ; and that my final offer :)

OpenStudy (amistre64):

\[b = 2s\pm2\]\[2s\pm2+3s=66\]\[5s\pm2=66\]\[5s=66\pm2\]\[s=\frac{66\pm2}{5}\]

OpenStudy (amistre64):

...just check their birth certificates :)

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

you are a horrible brother/sister if you dont know your own siblings' birthday and have to use algebra to figure out :(

OpenStudy (anonymous):

@superhelper As you see these can be tricky to set up. The answer given is fractional which isn't what was wanted, I think. I start by assuming that in 2 years Jack will be twice Jen's age in two years. I tried to solve the problem by guessing that Jen's age now is about 1/5 of 66. Jen = 11 now, will be 13; Jack is 24, will be 26; 3 times 11 + 24 = 57 Jen = 12 now, will be 14; Jack is 26, will be 28; 3 times 12 + 26 = 62 Jen = 13 now, will be 15; Jack is 28, will be 30; 3 times 13 + 28 = 67 This doesn't seem to give a whole number 66 answer! Change my assumption, in 2 years Jack will be twice Jen's age now. Jen = 12 now; Jack is 22, will be 24; 3 times 12 + 22 = 58 Jen = 13 now; Jack is 24, will be 26; 3 times 13 + 24 = 63 Jen = 14 now; Jack is 26, will be 28; 3 times 14 + 26 = 68 Seems to be something wrong with the problem… ? Recheck the text.

OpenStudy (anonymous):

the way i tried out this problem i didn't assume. what i wrote was right. i think you are suppose to come out with a decimal I think i just figured it out its hard to explain in words. but you do end up with a decimal

OpenStudy (anonymous):

Thanks for the update.

OpenStudy (anonymous):

yup